OpenStudy (anonymous):

Solve for 0 less than or equal to x less than 360. 4 sinx=-2

OpenStudy (campbell_st):

divide both sides by 4 sin(x) = -1/2 so the angles are in the 3rd and 4th quadrants. so find \[x = \sin^{-1}(\frac{1}{2})\] then the angle measures will be 3rd quadrant (180 + x) 4th quadrant (360 - x) hope this helps.

OpenStudy (anonymous):

yes it does! Thank you so much!